can you give me the breakdown of this math problem so that I can explain to a 11 year old to help with her homework
x + 1 = 1 1/4
3/4 - y = x

Answer by palanisamy(496)

Given, x + 1 = 1 1/4
x = 1 1/4 - 1
x = 1+1/4-1
x = 1/4
The second equation is
3/4-y = x
3/4+x = y
3/4+1/4 = y
(3+1)/4 = y
4/4 = y
1 = y
y = 1
Therefore the solution is x= 1/4 and y = 1